# A Microscopic Traffic Flow Model on Network with Destination-Aware V2V Communications and Rational Decision-Making

### v1

@@ Abstract
In this paper we carry out a computational study of a novel microscopic follow-the-leader model for traffic flow on road networks. We assume that each driver has its own origin and destination, and wants to complete its journey in minimal time. We also assume that each driver is able to take rational decisions at junctions and can change route while moving depending on the traffic conditions. The main novelty of the model is that vehicles can automatically and anonymously share information about their position, destination, and planned path when they are close to each other within a certain distance. The pieces of information acquired during the journey are used to optimize the route itself. In the limit case of a infinite communication range, we recover the classical Reactive User Equilibrium and Dynamic User Equilibrium.
@@ 1 Introduction
This paper is devoted to the study of a microscopic (agent-based) differential model for traffic flow on networks, where vehicles are able to communicate with each other within a certain distance. Note that we do not consider other than inter-vehicle communications, i.e. we assume that vehicles cannot send/receive information to/from the road infrastructure or any centralized server. We also assume that each vehicle has an assigned origin-destination (OD) pair, and all vehicles start moving from their origin at the same time. Moreover, each vehicle is equipped with an Onboard Processing Unit (OPU), which stores information about the geometry of the network, sends/receives information about position, destination, planned path, and is able to suggest rational decisions to the drivers regarding the minimum-time route towards destination. Since all vehicles are assumed to have such capabilities, we fall into a noncooperative game-theoretic approach for the overall traffic dynamics.
@@
Mathematical models for traffic flow on a single road have a very long story dating back to the 1950s, with a huge amount of research papers published on the topic. Strange to say, the mathematical and the engineering literature often proceeded in parallel without meeting, and arriving at the same conclusions from different points of view. Starting from the 1990s, the extensions to road networks became frequent, and now the literature covers extensively all possible approaches, including microscopic models, either differential (based on ordinary differential equations) and nondifferential (e.g., cellular automata), and macroscopic models, typically based on partial differential equations.
@@
  • Dynamic user equilibrium (DUE) Each driver knows other vehicles’ OD pairs, can predict the dynamics of all other vehicles, and is able to compute the fastest route on the basis of the current and future traffic conditions. Since this “predict & optimize” action is equally and independently taken by all drivers, we face a noncooperative game-theoretic problem. If a solution exists, it is a Nash-like (or Wardropian, in the traffic context) equilibrium, which corresponds to a scenario in which all drivers have chosen a route, and no driver can lower its time-to-destination unilaterally switching to a different route [25]. (An improvement is actually possible, but two or more drivers should agree to change their routes.) In a macroscopic setting, this scenario is also known as mean field game in the mathematical literature. Note also that optimization is not necessarily done with respect to the route choice (in other words, the behavior at junctions), but it can also be done with respect to the vehicle’s velocity [26, 27, 33]. Past drivers’ decisions can also be taken into account [6].
@@ 2 The models with full information
We also denote by $\ell$ the length of any vehicle, and by $\mathcal{L}$ their total length. It is important to note that there are two ways to vary the total number of cars $N_{\textsc{c}}$: one can do it keeping $\ell$ fixed or one can keep the relation $\ell=\frac{\mathcal{L}}{N_{\textsc{c}}}$ hold true, for a fixed $\mathcal{L}$. The first case, which we consider in this paper, it is used to truly change the degree of congestion of the network, while the second case is used to investigate possible micro-to-macro limit, since it translates the fact that the total length does not change when the number of vehicles tends to infinity, because the cars “shrink” accordingly.
@@ 2.1 Dynamic network loading
where $v_{\text{max}}$ is the maximal velocity. When a car reaches the end of its current road, it continues on the next road along its path as prescribed by the function nextroad.
One can note that this model does not guarantee that the distance between cars is larger than $\ell$ at any time, even if such constraint is satisfied at $t=0$. However, this is not a real issue because “smashed” cars can be found only in small intervals $[L_{\textsc{r}}-\ell,L_{\textsc{r}}]$ before any junction and their number is bounded by the number of incoming roads of that junction, see [15] for details.
@@ 2.2 Static route optimization via Dynamic Programming Principle
BB just requires to compute the shortest path for any car c from origin $\textsc{o}_{\textsc{c}}$ to destination $\textsc{d}_{\textsc{c}}$ once and for all, ignoring the presence of other cars on the network. To do that, one can simply employ the Dynamic Programming Principle (DPP).
@@
which acts as the cost (or weight) of the road r in the minimization. After that, for any car c, we define the value function $V_{\textsc{c}}(\textsc{j})$ which associates to any junction j the minimum time to reach the destination $\textsc{d}_{\textsc{c}}$ from that junction along the optimal path. Note that in this scenario the velocity is constant, then the shortest path coincides with the fastest path.
@@
$\left\{\begin{array}[]{l}V_{\textsc{c}}(\textsc{j})=\min\limits_{\textsc{r}\in\{r^{out}_{\textsc{j}}\}}\big{\{}V_{\textsc{c}}(\textsc{j}^{end}_{\textsc{r}})+w_{s}(\textsc{r})\big{\}},\qquad\forall\textsc{j}\neq\textsc{d}_{\textsc{c}},\\[11.38109pt] V_{\textsc{c}}(\textsc{d}_{\textsc{c}})=0.\end{array}\right.$ (3)
$V_{\textsc{c}}$ can be easily computed by a fixed-point algorithm, i.e. iterating the first equation in (3) until convergence, starting from $V_{\textsc{c}}(\textsc{j})=+\infty$ as initial guess for all $\textsc{j}\neq\textsc{d}_{\textsc{c}}$. Note that convergence is always obtained in a finite number of steps.
@@ 2.3 Reactive user equilibrium
    $w$ is the average travel time experienced by the cars passed along the road before time $t$, duly weighted in time to prioritize most recent vehicles.
@@
    $w$ is any increasing function of the number of vehicles on the road at time $t$.
@@
    $w$ is equal to $L_{\textsc{r}}/\hat{v}_{\textsc{r}}$, with $\hat{v}_{\textsc{r}}$ being the average velocity of the cars on the road r at time $t$.
In any case, it must be considered that one can get very large values for $w$ (or even +$\infty$+ ∞), because of the presence of queues. Moreover, yet another difficulty arises when DUE comes into play, we will detail this point in the next section.
@@ 2.4 Dynamic user equilibrium
The procedure for computing the DUE is different from the previous ones and more complicated, because the dynamic network loading (forward-in-time traffic flow model) and the route optimization (backward-in-time DPP) are fully coupled in the space-time and must be computed as one.
@@
is satisfied as well. Note that it is normal that, at equilibrium, $V_{\textsc{c}}=+\infty$ at some j, even for $t<T_{fin}$. This happens if there is not enough time to reach the destination before $T_{fin}$ starting from j at time $t$. The region in the space-time where $V_{\textsc{c}}$ is finite is called reachable set in control theory.
The actual computation of the solution (5) is not easy, and several methods were devised in the literature. We have used one of the most common, which consists in iterating the forward-in-time and the backward-in-time equations, keeping fixed $\textsc{nextroad}_{\textsc{c}}$ for all c in the first case and $w$ for all $t$ and r in the second case. Iterations keep going until convergence is reached. To add complexity, many possibilities exist to define convergence in this framework: one can check if all positions of all cars at all times stabilize at some iteration, or at least the functions $\textsc{nextroad}_{\textsc{c}}$ stabilize. We have preferred to employ an even weaker indicator, i.e. we check if the total travel time ttt, defined as the sum of times needed by all cars to complete their journey,
@@
eventually stabilizes. However, in any case the exact convergence is really difficult to obtain in a microscopic framework, and some initial conditions easily lead to instabilities, as reported in many papers; see, e.g., [37, 12]. To fix this, some techniques were proposed; among them, we have chosen the method of successive averages (MSA) (cf. also the fictitious play technique), consisting in using an average of the values of the weights $w$ obtained in previous iterations rather than just the last available value.
@@
Remark 2.1 (Perfect match between forward and backward dynamics).
@@
To test this hypothesis, we computed $w$ using a fourth method (denoted by M4 hereafter) in addition to the three methods M1, M2, M3 mentioned in Section 2.3. This method consists in running a fictitious simulation, which is an exact copy of the real one (same network, cars, and dynamics) with $t$ as initial time, where a probe vehicle starts from the beginning of the road r and moves until it reaches the end of the same road. Time needed to cover road r gives $w(t,\textsc{r})$. This choice is perfectly in line with the DPP in the space-time (6), which states that if one moves along the optimal path (unknown), “the time needed to reach the destination from any junction j at time $t$ equals the time $\delta t$ to reach the following junction $\textsc{j}^{*}$ plus the time to reach the destination starting from junction $\textsc{j}^{*}$ at time $t+\delta t$”. The time $\delta t$ is indeed the time computed in the fictitious simulation.
Actually, we observed that method M4 reduces instabilities for some initial conditions, but the improvement is not always consistent. Considering the major difficulty in implementing the method and the increased computational time, we think that this approach is not a good alternative to the other ones.
@@ 2.5.1 Test 0: Basic statistics
Many tests we will perform in the following will involve random origins and/or destinations. This is done because we want to keep an easy-to-get macroscale counterpart of our experiments, and random uniform initial and final conditions are associated to a constant initial and final densities on all over the network. Working with random variables requires to observe average quantities (expected values), therefore the question arises how many runs (repetitions) have to be performed in order to have meaningful statistics. The quantity which we will study the most will be the total travel time ttt defined in (7). Moreover, we expect RUE be the dynamics with larger deviations, so we run 500 simulations for that behavior, on the $5\times 5$ Manhattan-like network with random OD pairs, $L_{\textsc{r}}=50$ and $N_{\textsc{c}}=100$, and we plot the cumulative average
$\bar{X}_{r}:=\frac{1}{r}\sum_{i=1}^{r}\textsc{ttt}(i),\qquad r=1,\cdots,500,$
@@
Figure 2: Test 0. Cumulative average of ttt obtained for RUE, with $L_{\textsc{r}}=50$, $N_{\textsc{c}}=100$, $5\times 5$ Manhattan-like network with random OD pairs
@@
(|$\mathbb{P}\left(\text{\textbar}\bar{X}_{r}-\mu\text{\textbar}\leqslant\varepsilon_{r,\alpha}\right)=1-\alpha,\qquad\varepsilon_{r,\alpha}:=t_{\frac{\alpha}{2},r-1}\frac{\bar{\sigma}_{r}}{\sqrt{r}},$ (8)
where $\mu$ is the true mean of ttt (approximated by $\textsc{ttt}(500)$), $\bar{\sigma}_{r}$ is the standard deviation of the $r$ values, $t_{\frac{\alpha}{2},r-1}$ is the Student’s $t$ distribution with $r-1$ degrees of freedom, and $\alpha\in[0,1]$ is a parameter related to the reliability of the estimation. By setting $\varepsilon_{r,\alpha}=0.25$ and $\alpha=0.01$, we find that performing 300 runs is a reasonable choice for RUE, then it is for BB and DUE too. Consequently, in the tests that follow we will always provide results obtained by averaging over 300 runs.
@@
In this test we compare BB, RUE, and DUE on the two networks depicted in Fig. 1. Let us start with the simple 11-road network, which is specifically designed to highlight the differences among the three degrees of rationality. The destination is junction 4 for all $N_{\textsc{c}}=50$ cars, while initial positions are distributed across roads 2, 6, and 8. BB prescribes that all cars follow the shortest path, in particular cars starting from road 8 move straight through roads 0, 7, and finally 5 (Fig. 3a). RUE instead prescribes that cars starting from road 8 initially choose the shortest path, then they switch to the longer path 1-9-3-4 whenever the shortest becomes congested due to the merge with the cars coming from 2 and 6. When also the longer path becomes congested as well, they re-switch to the shortest one (Fig. 3b). Finally, DUE prescribes that cars starting form 8 are able to forecast the congestion on road 7 well before reaching the first junction, therefore they immediately start alternating between roads 0 and 1 in such a way that the shortest and the longer paths become competitive with each other (Fig. 3c).
@@
Figure 3: Test 1. Screenshots (at the same time) of the simulations on the 11-road simple network. They clearly show the difference among the three degrees of rationality

### v2

@@ Abstract
In this paper we carry out a computational study of a novel microscopic follow-the-leader model for traffic flow on road networks. We assume that each driver has its own origin and destination, and wants to complete its journey in minimal time. We also assume that each driver is able to take rational decisions at junctions and can change route while moving depending on the traffic conditions. The main novelty of the model is that vehicles can automatically and anonymously share information about their position, destination, and planned path when they are close to each other within a certain distance. The pieces of information acquired during the journey are used to optimize the route itself. In the limit case of an infinite communication range, we recover the classical Reactive User Equilibrium and Dynamic User Equilibrium.
@@ 1 Introduction
This paper is devoted to the study of a microscopic (agent-based) differential model for traffic flow on networks, where vehicles are able to communicate with each other within a certain distance. Note that we do not consider other than inter-vehicle communications, i.e. we assume that vehicles cannot send/receive information to/from the road infrastructure or any centralized server. We also assume that each vehicle has an assigned origin-destination (OD) pair, and all vehicles start moving from their origin at the same time. Moreover, each vehicle is equipped with an On-board Processing Unit (OPU), which stores information about the geometry of the network, sends/receives information about position, destination, planned path, and is able to suggest rational decisions to the drivers regarding the minimum-time route towards destination. Since all vehicles are assumed to have such capabilities, we fall into a noncooperative game-theoretic approach for the overall traffic dynamics.
@@
Mathematical models for traffic flow on a single road have a very long story dating back to the 1950s, with a huge amount of research papers published on the topic. It is useful to note that the mathematical and the engineering literature often proceeded in parallel without meeting, and arriving at the same conclusions from different points of view. Starting from the 1990s, the extensions to road networks became frequent, and now the literature covers extensively all possible approaches, including microscopic models, either differential (based on ordinary differential equations) and nondifferential (e.g., cellular automata), and macroscopic models, typically based on partial differential equations.
@@
  • Dynamic user equilibrium (DUE) Each driver knows other vehicles’ OD pairs, can predict the dynamics of all other vehicles, and is able to compute the fastest route on the basis of the current and future traffic conditions. Since this “predict & optimize” action is equally and independently taken by all drivers, we face a noncooperative game-theoretic problem. If a solution exists, it is a Nash-like (or Wardropian, in the traffic context) equilibrium, which corresponds to a scenario in which all drivers have chosen a route, and no driver can lower its time-to-destination unilaterally switching to a different route [4, 5, 6]. (An improvement is actually possible, but two or more drivers should agree to change their routes.) In a macroscopic setting, this scenario is also known as mean field game in the mathematical literature. Note also that optimization is not necessarily done with respect to the route choice (in other words, the behavior at junctions), but it can also be done with respect to the vehicle’s velocity [7, 8, 9]. Past drivers’ decisions can also be taken into account [10].
@@ 2 The models with full information
We also denote by $\ell$ the length of any vehicle, and by $\mathcal{L}$ their total length. It is important to note that there are two ways for varying the total number of cars $N_{\textsc{c}}$: one can do it keeping $\ell$ fixed or one can keep the relation $\ell=\frac{\mathcal{L}}{N_{\textsc{c}}}$ hold true, for a fixed $\mathcal{L}$. In the first setting (actually the one considered in this paper), one truly changes the number of vehicles adding mass to the system, and therefore increases the degree of congestion of the network; in the second setting, instead, one assumes that cars “shrink” as their number increases, preserving the total mass of the system. This way, we are just moving from a microscopic description to a macroscopic one, and, in the limit, cars disappear and are substituted by a density function with the same total mass [39].
@@ 2.1 Dynamic network loading
where $v_{\text{max}}$ is the maximal velocity. In this way, the velocity is proportional to the headway of the car c, and bounded in the interval $0,v_{\text{max}}]$. When a car reaches the end of its current road, it continues on the next road along its path as prescribed by the function nextroad.
As already noted in [39], the right-hand side in (1) is not continuous in time, since $\delta_{\textsc{c}}$ is not; Moreover, this model does not guarantee that the distance between cars is larger than $\ell$ at any time, even if such constraint is satisfied at $t=0$. However, pairs of cars closer than $\ell$ to each other can be found only in small intervals $L_{\textsc{r}}-\ell,L_{\textsc{r}}]$ just before any junction, and their number is bounded by the number of incoming roads of that junction, see [39, Sect. 3.1] for details. If this happens, the car behind stops and waits to have enough space ahead to move on.
@@ 2.2 Static route optimization via Dynamic Programming Principle
BB just requires to compute the shortest path for any car c from origin $\textsc{o}_{\textsc{c}}$ to destination $\textsc{d}_{\textsc{c}}$ once and for all, ignoring the presence of other cars on the network. Since vehicles’ dynamics is not involved here, this is a standard discrete optimization problem on a graph. To solve it, one can simply employ the Dynamic Programming Principle (DPP).
@@
which acts as the cost (or weight) of the road r in the minimization. After that, for any car c, we define the value function $V_{\textsc{c}}:\{1,\cdots,N_{\textsc{j}}\}\to\mathbb{R}^{+}$, which associates to any junction j the minimum time to reach the destination $\textsc{d}_{\textsc{c}}$ from that junction along the optimal path. Note that in this scenario the velocity is constant, then the shortest path coincides with the fastest path.
@@
$\left\{\begin{array}[]{l}V_{\textsc{c}}(\textsc{j})=\min\limits_{\textsc{r}\in\{r^{out}_{\textsc{j}}\}}\big\{V_{\textsc{c}}(\textsc{j}^{end}_{\textsc{r}})+w_{s}(\textsc{r})\big\},\qquad\forall\textsc{j}\neq\textsc{d}_{\textsc{c}},\\[11.38109pt] V_{\textsc{c}}(\textsc{d}_{\textsc{c}})=0,\end{array}\right$. (3)
where the set $\{r^{out}_{\textsc{j}}\}$, $j\neq\textsc{d}_{\textsc{c}}$, gives the set of admissible controls (in other words, the shortest path is found by choosing, in the optimal way, the outgoing roads from each junction, until $\textsc{d}_{\textsc{c}}$ is reached). $V_{\textsc{c}}$ can be easily computed by a fixed-point algorithm, i.e. iterating the first equation in (3) until convergence, starting from $V_{\textsc{c}}(\textsc{j})=+\infty$ as initial guess for all $\textsc{j}\neq\textsc{d}_{\textsc{c}}$. Note that convergence is always obtained in a finite number of steps.
@@ 2.3 Reactive user equilibrium
    $w(t,\textsc{r})$ is the average travel time experienced by the cars passed along the road r before time $t$, duly weighted in time to prioritize most recent vehicles.
@@
    $w(t,\textsc{r})$ is any increasing function of the number of vehicles on the road r at time $t$.
@@
    $w(t,\textsc{r})$ is equal to $L_{\textsc{r}}/\hat{v}_{\textsc{r}}$, with $\hat{v}_{\textsc{r}}$ being the average velocity of the cars on the road r at time $t$.
  4. M4.
    $w(t,\textsc{r})$ is the time needed to reach the end of the road r by a probe vehicle that starts moving at time $t$ from the beginning of the road r. The probe vehicle exists only in a fictitious simulation, which is an exact copy of the real one (same network, cars, and dynamics) and it is ran to the side of the main simulation, when needed.
In any case, it must be considered that one can get very large values for $w$ (or even $\infty$), because of the presence of queues. Moreover, yet another difficulty arises when DUE comes into play, we will detail this point in the next section.
@@ 2.4 Dynamic user equilibrium
The procedure for computing the DUE is different from the previous ones and more complicated, because the dynamic network loading (forward-in-time traffic flow model) and the route optimization (backward-in-time DPP) are fully coupled in the space-time and must be computed as one [43, 44, 4, 5, 6].
@@
is satisfied as well. The function $w$ is again computed by one of the methods M1-M4 mentioned above, and couples the forward and the backward parts of the system of equations. Note that it is normal that, at equilibrium, $V_{\textsc{c}}=+\infty$ at some j, even for $t<T_{fin}$. This happens if there is not enough time to reach the destination before $T_{fin}$ starting from j at time $t$. The region in the space-time where $V_{\textsc{c}}$ is finite is called reachable set in control theory [45, Chapt. IV, Sect. 1].
The actual computation of the solution (5) is not easy, and several methods were devised in the literature. We have used one of the most common, which consists in iterating the forward-in-time and the backward-in-time equations, keeping fixed $\textsc{nextroad}_{\textsc{c}}$ for all c in the first case and $w$ for all $t$ and r in the second case. Iterations keep going until convergence is reached, if any. To add complexity, many possibilities exist to define convergence in this framework: one can check if all positions of all cars at all times stabilize at some iteration, or at least the functions $\textsc{nextroad}_{\textsc{c}}$ stabilize. We have preferred to employ an even weaker indicator, i.e. we check if the total travel time ttt, defined as the sum of times needed by all cars to complete their journey,
@@
eventually stabilizes. Doing this, of course, there is the risk of masking different route distributions with the same global ttt. That said, in any case the exact convergence is really difficult to obtain in a microscopic framework, and some initial conditions easily lead to instabilities, as reported in many papers; see, e.g., [46, 47]. To fix this, some techniques were proposed; among them, we have chosen the method of successive averages (MSA) (cf. also the fictitious play technique), consisting in using an average of the values of the weights $w$ obtained in previous iterations rather than just the last available value.
@@
Remark 2.1 (Match between forward and backward dynamics).
@@
Method M1 is robust and temporally smooth but reacts slowly to rapid congestion changes, since it relies on past observations, therefore it may delay path updates. Moreover, it needs at least an additional parameter for prioritizing most recent vehicles.
Method M2 is the simplest and most economical one. However, it only reflects instantaneous density and does not distinguish different headway configurations, queues or stop-and-go patterns, making it rather coarse.
Method M3 is easy to compute, reacts quickly to current congestion, and is almost aligned with the microscopic dynamics, since speeds directly encode headway.
Finally, method M4 seems to guarantee perfect symmetry in forward and backward dynamics; also, it is in line with the DPP in the space-time (6), which states that if one moves along the optimal path (unknown), “the time needed to reach the destination from any junction j at time $t$ equals the time $\delta t$ to reach the following junction $\textsc{j}^{*}$ plus the time to reach the destination starting from junction $\textsc{j}^{*}$ at time $t+\delta t$”. The time $\delta t$ is indeed the time computed in the fictitious simulation. Actually, we observed that method M4 reduces instabilities for some initial conditions, but the improvement is not always consistent.
Considering the major difficulty in implementing M4, and the increased computational time, we think that the best balanced choice among all is M3; therefore, we adopt it in all our simulations.
@@ 2.5.1 Test 0: Basic statistics
Many tests we will perform in the following involve random origins and/or destinations. This is also done because we want to keep an easy-to-get macroscale counterpart of our experiments, and random uniform initial and final conditions are associated to a constant initial and final densities on all over the network. Working with random variables requires to observe average quantities (expected values), therefore the question arises how many runs (repetitions) have to be performed in order to have meaningful statistics. The quantity which we will study the most is the total travel time ttt defined in (7). Moreover, we expect RUE be the dynamics with larger deviations, so we run 300 simulations for that behavior on the Manhattan-like network with random OD pairs, $L_{\textsc{r}}=50$, and we plot the cumulative average
$\bar{X}_{r}:=\frac{1}{r}\sum_{i=1}^{r}\textsc{ttt}(i),\qquad r=1,\cdots,300,$
@@
(a) Comparison for $N_{\textsc{c}}=25$, $50$, and $100$ cars on $5\times 5$ Manhattan-like network
(b) Comparison for $3\times 3$, $5\times 5$, and $7\times 7$ Manhattan-like network with $N_{\textsc{c}}=100$ cars
Figure 2: Test 0. Cumulative average of ttt obtained for RUE, with $L_{\textsc{r}}=50$ and random OD pairs
@@
[…]
where $\mu$ is the true mean of ttt (approximated by $\textsc{ttt}(500)$), $\bar{\sigma}_{r}$ is the standard deviation of the $r$ values, $t_{\frac{\alpha}{2},r-1}$ is the Student’s $t$ distribution with r1r-1 degrees of freedom, and $\alpha\in[0,1]$ is a parameter related to the reliability of the estimation. By setting $\varepsilon_{r,\alpha}=0.25$ minutes and $\alpha=0.01$, we find that performing 300 runs is a reasonable choice for RUE, then it is for BB and DUE too. Consequently, in the tests that follow we will always provide results obtained by averaging over 300 runs.
@@
In this test we compare BB, RUE, and DUE on the two networks depicted in Fig. 1. Let us start with the simple 11-road network, which is specifically designed to highlight the differences among the three degrees of rationality. The destination is junction 4 for all $N_{\textsc{c}}=50$ cars, while initial positions are distributed across roads 2, 6, and 8. BB prescribes that all cars follow the shortest path, in particular cars starting from road 8 move straight through roads 0, 7, and finally 5 (Fig. 3a).111www.emilianocristiani.it/attach/paper˙trafficV2Vmicro˙smallnetwork˙50cars˙BB.mp4 RUE instead prescribes that cars starting from road 8 initially choose the shortest path, then they switch to the longer path 1-9-3-4 whenever the shortest becomes congested due to the merge with the cars coming from 2 and 6. When also the longer path becomes congested as well, they re-switch to the shortest one (Fig. 3b).222www.emilianocristiani.it/attach/paper˙trafficV2Vmicro˙smallnetwork˙50cars˙RUE.mp4 Finally, DUE prescribes that cars starting form 8 are able to forecast the congestion on road 7 well before reaching the first junction, therefore they immediately start alternating between roads 0 and 1 in such a way that the shortest and the longer paths become competitive with each other (Fig. 3c).333www.emilianocristiani.it/attach/paper˙trafficV2Vmicro˙smallnetwork˙50cars˙DUE.mp4
@@
Figure 3: Test 1. Three screenshots, corresponding to the same time $t$, of the simulation on the 11-road simple network (colors are just pictorial). They clearly show the difference among the three degrees of rationality